Solve for x; if log 7(x3+7)-log7(x+3)=2

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

LHS is log7 (x3+7)/(x+3)=log7(49)

x3+7/x+3=49

Now calculate the value of x